OpenStudy (anonymous):

find the equations for the asymptotes of the hyperbola (x^2)/36 - (y^2)/5 = 1

OpenStudy (anonymous):

y=?

OpenStudy (anonymous):

@Calcmathlete @waterineyes @zepp to the rescue..

OpenStudy (zepp):

First of all, tell me if your graph look like or and why :)

OpenStudy (anonymous):

First find the slope of the asymptotes. \(\huge m = ±\frac{rise}{run} = \frac{y}{x} = \frac{\sqrt{5}}{6}\) Since there is no translation for the center, the equation is \(\huge y = ±\frac{\sqrt{5}}{6}x\)

OpenStudy (zepp):

lol, answer farming